# Minkowski tensors for voxelized data: robust, asymptotically unbiased estimators

### v1

@@ 4.4 Convex test cases
In this section we present some simulation results showing how the algorithm works for specific simulated test data. We tested the algorithm for rectangles in dimension $d\in\{2,3\}$ as well as for a rectangle in dimension 2 with rounded vertices. To generate the data we fixed a grid $a\cdot{\mathbb{Z}}^{d}$ and intersected it with the rectangle. Table 4 shows formulas for the Minkowski tensors with rank at most 2 of a rectangle of the form $[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$. These formulas can be obtained by the following simplification of the Minkowski tensors of polytopes. For any polytope $P\subset{\mathbb{R}}^{d}$ we have
@@
$\Phi_{d}^{r,0}(P)$ $=\frac{1}{r!}\int_{P}x^{r}\,\mathrm{d}x,\qquad\qquad\qquad\Phi_{d}^{r,s}(P)=0,\qquad s>0,$
where $\mathcal{F}_{k}(P)$ is the set of the $k$-dimensional faces of the polytope $P$ and $N(P,F)$ is the normal cone of $P$ with respect to the face $F$. Recall that $\mathcal{H}^{k}$ denotes the $k$-dimensional Hausdorff measure and $\omega_{k}$ is the surface area of the unit sphere $\mathbb{S}^{k-1}$ in ${\mathbb{R}}^{k}$.
@@
Let $R=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$ be a two-dimensional rectangle with side lengths $a_{1}$ and $a_{2}$. Then the above formula yields for $r=2,s=0,k=1$ that
$\Phi_{1}^{2,0}(R)=\frac{1}{4}\sum_{F\in\mathcal{F}_{1}(R)}\int_{F}x^{2}\,\mathcal{H}^{1}(\mathrm{d}x).$
Using $x^{2}=x_{1}^{2}e_{1}^{2}+2x_{1}x_{2}e_{1}e_{2}+x_{2}^{2}e_{2}^{2}$, for an edge of the form $F=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times\{\pm\frac{a_{2}}{2}\}$ we obtain
$\int_{F}x^{2}\,\mathcal{H}^{1}(\mathrm{d}x)$ $=\int_{-\frac{a_{1}}{2}}^{\frac{a_{1}}{2}}y^{2}e_{1}^{2}\pm ya_{2}e_{1}e_{2}+\frac{a_{2}^{2}}{4}e_{2}^{2}\,\mathrm{d}y=\frac{a_{1}^{3}}{12}e_{1}^{2}+\frac{a_{1}a_{2}^{2}}{4}e_{2}^{2}.$
Therefore the summation over all the edges of $R$ yields
$\Phi_{1}^{2,0}(R)$ $=\frac{1}{24}\left((a_{1}^{3}+3a_{1}^{2}a_{2})e_{1}^{2}+(3a_{1}a_{2}^{2}+a_{2}^{3})e_{2}^{2}\right).$
@@
$\Phi_{1}^{1,1}(R)$ $=\frac{1}{2\pi}\sum_{F\in\mathcal{F}_{1}(R)}\int_{F}\int_{N(R,F)\cap\mathbb{S}^{1}}xu\,\mathcal{H}^{0}(\mathrm{d}u)\,\mathcal{H}^{1}(\mathrm{d}x).$
As before we start by computing the integral for an edge of the form $F=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times\{\pm\frac{a_{2}}{2}\}$. The normal cone of this set intersected with the unit sphere is given by {±$\{\pm e_{2}\}$. Since the tensor $xe_{2}$ can be written as $x_{1}e_{1}e_{2}+x_{2}e_{2}^{2}$, we obtain
$\int_{F}\int_{N(R,F)\cap\mathbb{S}^{1}}xu\,\mathcal{H}^{0}(\mathrm{d}u)\,\mathcal{H}^{1}(\mathrm{d}x)$ =±$=\pm\int_{F}xe_{2}\,\mathcal{H}^{1}(\mathrm{d}x)=\pm\int_{-\frac{a_{1}}{2}}^{\frac{a_{1}}{2}}ye_{1}e_{2}\pm\frac{a_{2}}{2}e_{2}^{2}\,\mathrm{d}y=\frac{a_{1}a_{2}}{2}e_{2}^{2}.$
@@
$\Phi_{1}^{1,1}(R)$ $=\frac{a_{1}a_{2}}{2\pi}(e_{1}^{2}+e_{2}^{2})=\frac{a_{1}a_{2}}{2\pi}\,Q.$
The results of the algorithm for a two-dimensional rectangle can be found in Appendix A and are very close to the exact values. We also tested a 3-dimensional rectangle in the same way and obtained very good results. Further we considered a 2-dimensional rectangle with “rounded corners” and the results again were very close to the exact values (relative deviations of less than 1% for rectangles of varying sizes with up to 1.3 million points). The detailed results, along with those for a 3-dimensional rectangle, where we once again achieved very good outcomes, can be found in Appendix A.
Table 4: Formulas for the Minkowski tensors with rank at most 2 of a rectangle of the form $[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$, where the values of $\Phi_{k}^{1,0}$, $\Phi_{k}^{0,1}$ vanish for all $k\in\{0,1,2\}$.
@@
$\Phi_{0}^{0,0}$ 1 $\Phi_{1}^{0,0}$ $a_{1}+a_{2}$ $\Phi_{2}^{0,0}$ $a_{1}\cdot a_{2}$
$\Phi_{0}^{0,2}$ $\frac{1}{4\pi}\left(e_{1}^{2}+e_{2}^{2}\right)$ $\Phi_{1}^{0,2}$ $\frac{1}{4\pi}\left(a_{2}e_{1}^{2}+a_{1}e_{2}^{2}\right)$ $\Phi_{2}^{0,2}$ 0
$\Phi_{0}^{2,0}$ $\frac{1}{8}\left(a_{1}^{2}e_{1}^{2}+a_{2}^{2}e_{2}^{2}\right)$ $\Phi_{1}^{2,0}$ $\frac{1}{24}\sum_{i=1}^{2}\left(a_{i}^{3}+3a_{i}^{2}a_{j}\right)e_{i}^{2}$ $\Phi_{2}^{2,0}$ $\frac{1}{24}\left(a_{1}^{3}a_{2}e_{1}^{2}+a_{2}^{3}a_{1}e_{2}^{2}\right)$
$\Phi_{0}^{1,1}$ $\frac{1}{2\pi}\left(a_{1}e_{1}^{2}+a_{2}e_{2}^{2}\right)$ $\Phi_{1}^{1,1}$ $\frac{a_{1}a_{2}}{2\pi}\left(e_{1}^{2}+e_{2}^{2}\right)$ $\Phi_{2}^{1,1}$ 0
@@ 4.5 Nonconvex test cases
If $A\subset{\mathbb{R}}^{d}$ is a set of positive reach, the boundary $\partial A$ of $A$ is of class $C^{2}$, and $A$ has a unique exterior unit outer normal vector $\nu(A,x)$ at each $x\in\partial A$, then
$\Lambda_{m}(A,\cdot)=\frac{1}{\omega_{d-m}}\int_{\partial A}\mathds{1}\{(x,\nu(A,x))\in\cdot\}\sum_{|I|=d-1-m}\prod_{i\in I}k_{i}(A,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}),$
where $k_{1}(A,x),\ldots,k_{d-1}(A,x)\in{\mathbb{R}}$ are the principal curvatures of $A$ at $x\in\partial A$ and the summation extends over all subsets $I\subset\{1,\ldots,d-1\}$ of cardinality $|I|=d-1-m$ (if $m=d-1$ the empty product is interpreted as $1$). Thus we obtain, for $k\in\{0,\dots,d-1\}$,
$\Phi^{r,s}_{k}(A)=\frac{1}{r!s!\omega_{d-k+s}}\int_{\partial A}x^{r}\nu(A,x)^{s}\sum_{|I|=d-1-m}\prod_{i\in I}k_{i}(A,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}).$
@@
Let $A=\{x\in{\mathbb{R}}^{d}:\rho_{1}\leq\|x\|\leq\rho_{2}\}$ denote a spherical shell with inner radius $\rho_{1}$ and outer radius $\rho_{2}$, where $0\leq\rho_{1}<\rho_{2}<\infty$. Observe that the reach of $A$ is $\rho_{1}$ whenever $\rho_{1}>0$ (if $\rho_{1}=0$, then $A$ is a Euclidean ball with radius $\rho_{2}$). Since the principal curvatures of $A$ at $x\in\partial A$ are equal to $\rho_{2}^{-1}$, if $\|x\|=\rho_{2}$, and equal to $-\rho_{1}^{-1}$, if $\|x\|=\rho_{1}>0$, we simply write $k(A,x)$ for these principal curvatures and get
@@
where the transformation formula and [27, Lemma 7] have been used to determine the integral.
If $k=d-1$, this specializes to
@@
Table 5 shows simulation results for the 2-dimensional spherical shell with inner radius $\rho_{1}=1$ and outer radius $\rho_{2}=2$.
Table 5: Results for the 2-dimensional spherical shell with inner radius $\rho_{1}=1$ and outer radius $\rho_{2}=2$ intersected with a grid of resolution $a=0.005$. The parameter choices were $n=50$ and $R_{n}=1.8$ (0.9 times the reach of the spherical shell). We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Values Algorithm Tensors Values Algorithm
[…]
@@
Example 4.11.
Let $R_{a}=[-\frac{a_{1}}{2},\frac{a_{1}}{2}]\times[-\frac{a_{2}}{2},\frac{a_{2}}{2}]\subset{\mathbb{R}}^{2}$ be a two-dimensional rectangle with side lengths $a_{1}$ and $a_{2}$ and $a=(a_{1},a_{2})^{\top}\in{\mathbb{R}}^{2}$. For $0<a_{1}<b_{1}$ and $0<a_{2}<b_{2}$ we consider $R_{a,b}:=R_{b}\setminus R_{a}^{\circ}$, i.e., the interior of the rectangle $R_{a}$ is removed from the enclosing rectangle $R_{b}$. The set $R_{a,b}$ does not have positive reach, but it is a polyconvex set and in particular a finite union of compact sets with positive reach. The surface tensors $\Phi_{1}^{r,s}(R_{a,b})$ can be expressed in terms of the surfaces tensors of $R_{a}$ and $R_{b}$ by
$\Phi_{1}^{r,s}(R_{a,b})=\Phi_{1}^{r,s}(R_{b})+(-1)^{s}\Phi_{1}^{r,s}(R_{a}),$
where $r,s\in{\mathbb{N}}_{0}$; see Table 4 for explicit formulas. Table 6 shows simulation results for $R_{a,b}$ with $a=(1,2)^{\top}$, $b=(3,5)^{\top}$.
Table 6: Results for $R_{a,b}$ with $a=(1,2)^{\top}$, $b=(3,5)^{\top}$ intersected with a grid of resolution $a=0.01$. The parameter choices were $n=50$ and $R_{n}=0.45$ ($0.9\cdot 0.5$ times the length of the smaller side of $R_{a}$). We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Values Algorithm Tensors Values Algorithm

### v2

@@ 5.1 Convex test cases
In this section we present some simulation results showing how the Voronoi-LSQ algorithm works for specific simulated test data. We tested the algorithm for rectangles in dimension $d\in\{2,3\}$ as well as for parallel sets of a rectangle in dimension 2. Table 4 shows formulas for the Minkowski tensors with rank at most 2 of a rectangle of the form $\frac{p_{1}}{2},\frac{p_{1}}{2}]\times[-\frac{p_{2}}{2},\frac{p_{2}}{2}]\subset{\mathbb{R}}^{2}$. These formulas can be obtained by the following simplification of the Minkowski tensors of polytopes. For a polytope $P\subset{\mathbb{R}}^{d}$ we obtain from (2.4), (2.6) and (2.7) that
@@
$\Phi_{d}^{r,0}(P)$ $=\frac{1}{r!}\int_{P}x^{r}\,\mathrm{d}x,\qquad\qquad\qquad\Phi_{d}^{r,s}(P)=0,\qquad s>0.$
@@
Let $R_{p}=[-\frac{p_{1}}{2},\frac{p_{1}}{2}]\times[-\frac{p_{2}}{2},\frac{p_{2}}{2}]\subset{\mathbb{R}}^{2}$ be a two-dimensional rectangle with side lengths $p_{1}$ and $p_{2}$ and $p=(p_{1},p_{2})^{\top}\in{\mathbb{R}}^{2}$. Then the above formula yields for $r=2,s=0,k=1$ that
$\Phi_{1}^{2,0}(R_{p})=\frac{1}{4}\sum_{F\in\mathcal{F}_{1}(R_{p})}\int_{F}x^{2}\,\mathcal{H}^{1}(\mathrm{d}x).$
Using $x^{2}=x_{1}^{2}e_{1}^{2}+2x_{1}x_{2}e_{1}e_{2}+x_{2}^{2}e_{2}^{2}$, for an edge of the form $F=[-\frac{p_{1}}{2},\frac{p_{1}}{2}]\times\{\pm\frac{p_{2}}{2}\}$ we obtain
$\int_{F}x^{2}\,\mathcal{H}^{1}(\mathrm{d}x)$ $=\int_{-\frac{p_{1}}{2}}^{\frac{p_{1}}{2}}y^{2}e_{1}^{2}\pm yp_{2}e_{1}e_{2}+\frac{p_{2}^{2}}{4}e_{2}^{2}\,\mathrm{d}y=\frac{p_{1}^{3}}{12}e_{1}^{2}+\frac{p_{1}p_{2}^{2}}{4}e_{2}^{2}.$
Therefore the summation over all the edges of $R_{p}$ yields
$\Phi_{1}^{2,0}(R_{p})$ $=\frac{1}{24}\left((p_{1}^{3}+3p_{1}^{2}p_{2})e_{1}^{2}+(3p_{1}p_{2}^{2}+p_{2}^{3})e_{2}^{2}\right).$
@@
$\Phi_{1}^{1,1}(R_{p})$ $=\frac{1}{2\pi}\sum_{F\in\mathcal{F}_{1}(R_{p})}\int_{F}\int_{N(R_{p},F)\cap\mathbb{S}^{1}}xu\,\mathcal{H}^{0}(\mathrm{d}u)\,\mathcal{H}^{1}(\mathrm{d}x).$
As before we start by computing the integral for an edge of the form $F=[-\frac{p_{1}}{2},\frac{p_{1}}{2}]\times\{\pm\frac{p_{2}}{2}\}$. The normal cone of this set intersected with the unit sphere is given by $\{\pm e_{2}\}$. Since the tensor $xe_{2}$ can be written as $x_{1}e_{1}e_{2}+x_{2}e_{2}^{2}$, we obtain
$\int_{F}\int_{N(R_{p},F)\cap\mathbb{S}^{1}}xu\,\mathcal{H}^{0}(\mathrm{d}u)\,\mathcal{H}^{1}(\mathrm{d}x)$ $=\pm\int_{F}xe_{2}\,\mathcal{H}^{1}(\mathrm{d}x)=\pm\int_{-\frac{p_{1}}{2}}^{\frac{p_{1}}{2}}ye_{1}e_{2}\pm\frac{p_{2}}{2}e_{2}^{2}\,\mathrm{d}y=\frac{p_{1}p_{2}}{2}e_{2}^{2}.$
@@
$\Phi_{1}^{1,1}(R_{p})$ $=\frac{p_{1}p_{2}}{2\pi}(e_{1}^{2}+e_{2}^{2})=\frac{p_{1}p_{2}}{2\pi}\,Q.$
The results of the Voronoi-LSQ algorithm for a two-dimensional rectangle can be found in the supplement and are very close to the exact values. We also tested a three-dimensional rectangle in the same way and obtained very good results. Further we considered parallel sets of a two-dimensional rectangle and the results again were very close to the exact values (relative deviations of less than 1% for rectangles of varying sizes with up to 1.3 million points). The detailed results, along with those for a three-dimensional rectangle, where we once again achieved very good outcomes, can be found in the supplementary materials.
Table 4: Formulas for the Minkowski tensors with rank at most 2 of a rectangle of the form $\frac{p_{1}}{2},\frac{p_{1}}{2}]\times[-\frac{p_{2}}{2},\frac{p_{2}}{2}]\subset{\mathbb{R}}^{2}$, where the values of $\Phi_{k}^{1,0}$, $\Phi_{k}^{0,1}$ vanish for all $k\in\{0,1,2\}$.
@@
$\Phi_{0}^{0,0}$ 1 $\Phi_{1}^{0,0}$ $p_{1}+p_{2}$ $\Phi_{2}^{0,0}$ $p_{1}\cdot p_{2}$
$\Phi_{0}^{0,2}$ $\frac{1}{4\pi}\left(e_{1}^{2}+e_{2}^{2}\right)$ $\Phi_{1}^{0,2}$ $\frac{1}{4\pi}\left(p_{2}e_{1}^{2}+p_{1}e_{2}^{2}\right)$ $\Phi_{2}^{0,2}$ 0
$\Phi_{0}^{2,0}$ $\frac{1}{8}\left(p_{1}^{2}e_{1}^{2}+p_{2}^{2}e_{2}^{2}\right)$ $\Phi_{1}^{2,0}$ $\frac{1}{24}\sum_{i=1}^{2}\left(p_{i}^{3}+3p_{i}^{2}p_{j}\right)e_{i}^{2}$ $\Phi_{2}^{2,0}$ $\frac{1}{24}\left(p_{1}^{3}p_{2}e_{1}^{2}+p_{2}^{3}p_{1}e_{2}^{2}\right)$
$\Phi_{0}^{1,1}$ $\frac{1}{2\pi}\left(p_{1}e_{1}^{2}+p_{2}e_{2}^{2}\right)$ $\Phi_{1}^{1,1}$ $\frac{p_{1}p_{2}}{2\pi}\left(e_{1}^{2}+e_{2}^{2}\right)$ $\Phi_{2}^{1,1}$ 0
@@ 5.2 Nonconvex test cases
If $A\subset{\mathbb{R}}^{d}$ is a set of positive reach which is the closure of its interior, the boundary $\partial A$ of $A$ is of class $C^{2}$, and $A$ has a unique exterior unit outer normal vector $\nu(A,x)$ at each $x\in\partial A$ (that is, $\partial A=\partial^{1}A$), then we obtain as in (2.5) that
$\Lambda_{j}(A,\cdot)=\frac{1}{\omega_{d-j}}\int_{\partial A}\mathds{1}\{(x,\nu(A,x))\in\cdot\}\sum_{|I|=d-1-j}\prod_{i\in I}k_{i}(A,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}),$
where $k_{1}(A,x),\ldots,k_{d-1}(A,x)\in{\mathbb{R}}$ are the principal curvatures of $A$ at $x\in\partial A$ (with respect to the outer unit normal) and the summation extends over all subsets $I\subseteq\{1,\ldots,d-1\}$ of cardinality |I|=d1j|I|=d-1-j (if j=d1j=d-1 the empty product is interpreted as $1$). Note that if $A$ is convex, then $k_{i}(K,x)\geq 0$ for $i=0,\ldots,d-1$. Thus, for $k\in\{0,\dots,d-1\}$, it follows from (2.6) that
$\Phi^{r,s}_{k}(A)=\frac{1}{r!s!\omega_{d-k+s}}\int_{\partial A}x^{r}\nu(A,x)^{s}\sum_{|I|=d-1-k}\prod_{i\in I}k_{i}(A,x)\,\mathcal{H}^{d-1}(\mathrm{d}{x}).$
@@
Let $A=\{x\in{\mathbb{R}}^{d}:\rho_{1}\leq\|x\|\leq\rho_{2}\}$ denote a spherical shell with inner radius $\rho_{1}$ and outer radius $\rho_{2}$, where $0\leq\rho_{1}<\rho_{2}<\infty$. Observe that the reach of $A$ is $\rho_{1}$ whenever $\rho_{1}>0$ (if $\rho_{1}=0$, then $A$ is a Euclidean ball with radius $\rho_{2}$). Since the principal curvatures of $A$ at $x\in\partial A$ are equal to $\rho_{2}^{-1}$, if $\|x\|=\rho_{2}$, and equal to $\rho_{1}^{-1}$, if $\|x\|=\rho_{1}>0$, we simply write $k(A,x)$ for these principal curvatures and get
@@
where the transformation formula and [31, Lemma 7] have been used to determine the integral.
If k=d1k=d-1, this specializes to
@@
Table 5 shows simulation results of the Voronoi-LSQ algorithm for the two-dimensional spherical shell with inner radius $\rho_{1}=1$ and outer radius $\rho_{2}=2$.
Table 5: Results of the Voronoi-LSQ algorithm for the two-dimensional spherical shell with inner radius $\rho_{1}=1$ and outer radius $\rho_{2}=2$ intersected with a grid of resolution $\Delta=0.005$. The parameter choices were $n=50$ and $R_{n}=0.9$ (0.9 times the reach of the spherical shell). We took the average of 25 renditions. All values are rounded to the fourth significant digit.
Tensors Exact value Voronoi-LSQ Tensors Exact value Voronoi-LSQ
$\Phi_{0}^{0,0}$ 0 $10^{-3}$ $\Phi_{1}^{0,0}$ 9.4248 9.4441
@@
Table 6: Results of the Voronoi-LSQ algorithm for $R_{p,q}$ with $p=(1,2)^{\top}$, $q=(3,5)^{\top}$ intersected with a grid of resolution $\Delta=0.01$. The parameter choices were $n=50$ and $R_{n}=0.45$ ($0.9\cdot 0.5$ times the length of the smaller side of $R_{p}$). We took the average of 25 renditions.
Tensors Exact value Voronoi-LSQ Tensors Exact value Voronoi-LSQ
@@
Example 5.3.
Let $R_{p}=[-\frac{p_{1}}{2},\frac{p_{1}}{2}]\times[-\frac{p_{2}}{2},\frac{p_{2}}{2}]\subset{\mathbb{R}}^{2}$ be a two-dimensional rectangle with side lengths $p_{1}$ and $p_{2}$ and $p=(p_{1},p_{2})^{\top}\in{\mathbb{R}}^{2}$. For $0<p_{1}<q_{1}$ and $0<p_{2}<q_{2}$ we consider $R_{p,q}:=R_{q}\setminus R_{p}^{\circ}$, i.e., the interior of the rectangle $R_{p}$ is removed from the enclosing rectangle $R_{q}$. The set $R_{p,q}$ does not have positive reach, but it is a polyconvex set and in particular a finite union of compact sets with positive reach. The surface tensors $\Phi_{1}^{r,s}(R_{p,q})$ can be expressed in terms of the surfaces tensors of $R_{p}$ and $R_{q}$ by
$\Phi_{1}^{r,s}(R_{p,q})=\Phi_{1}^{r,s}(R_{q})+(-1)^{s}\Phi_{1}^{r,s}(R_{p}),$
where $r,s\in{\mathbb{N}}_{0}$; see Table 4 for explicit formulas. Table 6 shows simulation results of the Voronoi-LSQ algorithm for $R_{p,q}$ with $p=(1,2)^{\top}$, $q=(3,5)^{\top}$.
5.3 Convergence analysis
Figure 7: The relative error in estimating $\Phi_{1}^{0,0}$ (which equals half of the surface area) of the sets from Examples 5.1, 5.2 and 5.3 using the Voronoi-LSQ algorithm. The test data were obtained by intersecting the sets with a grid of resolution $\Delta$ and the relative error is shown as a function of $\Delta$.
After comparing the results of the Voronoi-LSQ algorithm for a fixed choice of $R_{n}$ with the true values, we now investigate the dependence of the error on the resolution of the test sets. Recall that the test data were generated by intersecting the test sets with a grid of a given resolution. Figure 7 shows the relative error of the estimated surface areas for the sets from Examples 5.1, 5.2 and 5.3, plotted as a function of the resolution, which ranges over about three orders of magnitude. Over this range, the error in the surface area drops by about five orders of magnitude. All algorithm parameters and the precise specifications of the example sets were adopted exactly as in the simulations presented in Subsections 5.1 and 5.2. As theoretically expected, the error of the algorithm’s results converges to zero as the resolution decreases for the rectangle and the circular shell. Our simulations suggest that this holds for the rectangular shell as well, although no theoretical proof is available. A relative error of 0.1% is achieved at a resolution of 0.002 (circular shell), 0.01 (rectangular shell), and 0.05 (rectangle), respectively.